OpenStudy (anonymous):

I am adding the two rational expressions (x/x+1)+(x/x-1). This is what I have so far. First you multiply the numerator and denominator by the denominator of the other fraction. (x-1) x/(x-1)(x+1)+ x/(x-1) = (x²-x)/(x²+x-x-1) Then you must combine like terms in the denominator. (x²-x)/(x²-1) I'm kind of stuck. Am I supposed to factor the expression?

OpenStudy (anonymous):

\[\frac{ x }{ x+1 }+\frac{ x }{ x-1 }=x \left[ \frac{ 1 }{ x+1 }+\frac{ 1 }{ x-1 } \right]=x \frac{ x-1+x+1 }{ \left( x+1 \right)\left( x-1 \right) }=?\]

OpenStudy (anonymous):

Is it supposed to be 2x/(x^2-1) @surjithayer

OpenStudy (anonymous):

it is \[\frac{ 2 x^2 }{ x^2-1 }\]